### v5

@@ 2 Targeting
It is important to note here that the values $U_{z}(t)$ and therefore the targeting maps $Z^{\ast}$ and $T^{\ast}$ are not observable; any assumption on targeting instruments and targeted treatments must be a priori and context-dependent. As we will see, these prior assumptions sometimes have consequences that can be tested.
@@
Under one-to-one targeting, we will often write “$z=t$” if $z$ targets $t$; this is without loss of generality. Let us illustrate these varieties of targeting on Example 2.
@@
2.1 Consequences of One-to-One Targeting
In this subsection, we impose
Assumption 4 (One-to-one Targeting).
Targeting is one-to-one.
Remember that under Assumption 4, we can relabel instrument values so that if $t$ is targeted, then it is targeted by $z=t$. Moreover, $t^{\ast}(z)$ must equal $z$.
This implies some useful restrictions on response-groups.
@@ 2.1 Consequences of One-to-One Targeting
Under Assumptions 1, 2, and 4, take a targeted treatment $t\in\mathcal{T}^{\ast}$.
@@
    If an observation $i$ has $T_{i}(t)=0$, then it never receives treatment $t$: $T_{i}(z)\neq t$ for all $z\in\mathcal{Z}$.
@@
    As a consequence, all response-groups $C_{R}$ with $R(t)=0$ and $R(z)=t$ for some $z\neq t$ are empty.
@@
This is easily rephrased in terms of the response-vectors of definition 1. First note that in the $3\times 3$ case, there are a priori $27\displaystyle 3^{3}=$ 27 response-vectors, $R=000$ to $R=222$, with corresponding response-groups $C_{000}$ to $C_{222}$. Groups $C_{ddd}$ are “always-takers” [^14] Observations in group $C_{000}$ are usually called the “never-takers”. We prefer not to break the symmetry in our notation. We hope this will not cause confusion. of treatment value $d$. All other groups are “compliers” of some kind, in that their treatment changes under some changes in the instrument. We will also pay special attention to some non-elemental groups. For instance, $C_{0\ast 2}$ will denote the group who is assigned treatment 0 under $z=0$ and treatment 2 under $z=2$, and any treatment under $z=1$. That is,
@@
Assumptions 2 and 4 together imply the emptiness of four composite groups out of the 27 possible. For any treatment value $\tau$, Proposition 1(ii) rules out group $C_{10\tau}$ since this group has $R(1)=0$ and $R(0)=1$. It rules out $C_{\tau 01}$ as $R(1)=0$ and $R(2)=1$. This eliminates the composite groups $C_{10\ast}$ and $C_{\ast 01}$. The same argument applies to composite groups $C_{\ast 20}$ and $C_{2\ast 0}$, which have $R(2)=0$ and $R(1)=2$ or $R(2)=2$.
These four composite groups correspond to 10 elemental groups [^15] Specifically, they are: $C_{100},C_{101},C_{102},C_{001},C_{201},C_{020},C_{120},C_{220},C_{200}$, and $C_{210}$.. This still leaves us with 17 elemental groups, and potentially complex assignment patterns. Consider for instance Figure 2. It shows one possible configuration for the $3\times 3$ model; the positions for $P_{0},P_{1}$ and $P_{2}$ are consistent with Assumptions 2 and 4.
[…]
@@
Heckman and Pinto (2018, pp. 16–20), Pinto (2021), and Kirkeboen, Leuven, and Mogstad (2016) also studied the $3\times 3$ model; they proposed sets of assumptions that identify some treatment effects. The example in Heckman and Pinto (2018, pp. 16–20) is rather specific. We show in Appendix E how to apply our framework to the Moving to Opportunity experiment studied in Pinto (2021). The setup in Kirkeboen, Leuven, and Mogstad (2016) is most similar to ours; we will return to the differences between our approach and theirs in Section 3.3.
@@ 2.2 Strict Targeting
Suppose for instance the data comes from a randomized experiment, where the instrument value $z=t$ targets treatment $t$. If compliance is imperfect, an individual will trade off the benefits from switching to a treatment $t^{\prime}\neq t$ with the costs of the effort required. Strict targeting obtains when the cost of switching to $t^{\prime}$ do not depend on the value of $t$.
@@
2.3 Consequences of Strict Targeting
@@
Assumption 5 (Strict targeting).
@@
Under strict targeting, the values of $U_{z}(t)$ are given in Table 2.
@@
$z\not\in Z^{\ast}$ $\underline{U}(t)$
Table 2: Values of $U_{z}(t)$ under strict targeting
Consider an observation $i$ under strict targeting. If it is assigned an instrument value $z$, it can end up with one of the treatment values $t$ that $z$ targets (if any), with a value $\bar{U}(t)+u_{it}$ in the ARUM. Alternatively, if its treatment is some $t^{\prime}$ that $z$ does not target, then the ARUM value will be $\underline{U}(t^{\prime})+u_{it^{\prime}}$. This motivates the following definition.
Definition 5 (Top targeted and top alternative treatments).
Take any observation $i$ in the population.
@@
    For any targeting instrument $z\in\mathcal{Z}^{\ast}$, let
    $V^{\ast}_{i}(z)=\max_{t\in T^{\ast}(z)}(\bar{U}(t)+u_{it})$
    and $T^{\ast}_{i}(z)\subset T^{\ast}(z)$ denote the set of maximizers. We call the elements of $T^{\ast}_{i}(z)$ the top targeted treatments for observation $i$ under instrument value $z$.
@@
    Also define
    $\underline{V}_{i}=\max_{t\in\mathcal{T}}\left(\underline{U}(t)+u_{it}\right)$
    and let $\underline{T}_{i}\subset\mathcal{T}$ denote the set of maximizers. We call the elements of $\underline{T}_{i}$ the top alternative treatments for observation $i$.
@@
    The sets $T^{\ast}_{i}(z)$ and $\underline{T}_{i}$ are singletons [^16] Note that this follows from our assumption that the distribution of the random vector ${(u_{it})}_{t\in\mathcal{T}}$ is absolutely continuous; however, it does not extend to the sets $Z^{\ast}(t)$ and $T^{\ast}(z)$, which can still have several elements. with probability 1; we let $t^{\ast}_{i}(z)$ and $\underline{t}_{i}$ denote the top targeted treatment and the top alternative treatment.
The term “top alternative treatment” may read like a misnomer since the maximization runs over all treatment values. The following result justifies it; more importantly, it shows that strict targeting imposes a lot of structure on the mapping from instruments to treatments.
Proposition 2 (Response groups under strict targeting).
Let Assumptions 1, 2, 3, and 5 hold. Let $i$ be any observation in the population. For any instrument value $z$, $T_{i}(z)$ is either the top targeted treatment or the top alternative treatment. If $z$ is not a targeting instrument, $T_{i}(z)$ can only be the top alternative treatment. That is:
  1. (i)
    if $z\in\mathcal{Z}^{\ast}$, then $T_{i}(z)$ is $t^{\ast}_{i}(z)$ if $V^{\ast}_{i}(z)>\underline{V}_{i}$; if $V^{\ast}_{i}(z)<\underline{V}_{i}$, then $T_{i}(z)=\underline{t}_{i}$ and $\underline{t}_{i}$ is not targeted by $z$.
  2. (ii)
    if $z\not\in\mathcal{Z}^{\ast}$, then $T_{i}(z)$ is $\underline{t}_{i}$.
Note that in a sense, all instrument values in $\mathcal{Z}\,\setminus\,\mathcal{Z}^{\ast}$ are equivalent under strict targeting. If $z$ and $z^{\prime}$ are both in $\mathcal{Z}^{\ast}$, then the functions $U_{z}$ and $U_{z^{\prime}}$ coincide on all of $\mathcal{T}$ and the counterfactual treatments $T_{i}(z)$ and $T_{i}(z^{\prime})$ must be in $\underline{T}_{i}$ for any observation $i$.
In the $2\times 2$ model, we have $\bar{U}(1)=U_{1}(1)$ and $\underline{U}(0)=\underline{U}(1)=0$. A complier is an observation $i\in C_{01}$; it is in treatment arm $t=0$ when $z=0$ and in $t^{\ast}(1)=1$ when $z=1$. In our more general model, it seems natural to define a $t$-complier as an observation $i$ that is in treatment arm $t$ when assigned an instrument value $z$ such that $t^{\ast}_{i}(z)=t$, and only then. This is, clearly, a composite group. Take the $3\times 3$ model as an example, and assume that $t^{\ast}(1)=1$. Then the set of 11 11-compliers consists of the five response-groups $C_{010}$, $C_{012}$, $C_{111}$, $C_{112}$, and $C_{212}$.
2.4 Strict one-to-one targeting
We now impose one-to-one targeting (Assumption 4) as well as strict targeting. Under one-to-one targeting, the sets $Z^{\ast}(t)$ and $T^{\ast}(z)$ are singletons; and each targeting instrument $z$ can be relabeled as the treatment value $t=t^{\ast}_{i}(z)$ that it targets.
Corollary 1 (Treatment assignment under strict, one-to-one targeting).
Take any observation $i$. Let $A_{i}$ be the (possibly empty) subset of $t\in\mathcal{T}^{\ast}$ such that $T_{i}(t)=t$. Then under Assumptions 1 to 5,
  1. 1.
    $T_{i}(t)=\underline{t}_{i}$ for all $t\in\mathcal{T}\,\setminus\,A_{i}$;
  2. 2.
    if $\underline{t}_{i}$ is a targeted treatment, it must belong to $A_{i}$.
The pair $(A_{i},\underline{t}_{i})$ defines an elemental response group which we denote $C(A_{i},\underline{t}_{i})$. The family of sets $\{C(A,t)\;|\;A\subset\mathcal{T}^{\ast},t\not\in\mathcal{T}^{\ast}\,\setminus\,A\}$ form a partition of the set of observations.
@@ 2.4 Strict one-to-one targeting
If $\underline{t}_{i}\not\in\mathcal{T}^{\ast}$ and the set $A_{i}$ is non-empty, then the observation $i$ is what one could call a strict $A_{i}$-complier: when the value of the instrument moves from $\mathcal{Z}\,\setminus\,A_{i}$ to $t\in A_{i}$, observation $i$ switches from its top alternative treatment $\underline{t}_{i}$ to the treatment $t$. In the 3-by-3 model with $\mathcal{T}^{\ast}=\{1,2\}$, there are three groups of strict compliers: $C_{010}=C(\{1\},0)$, $C_{002}=C(\{2\},0)$, and $C_{012}=C(\{1,2\},0)$.
Strict one-to-one targeting brings us very close to the main identifying assumption in Heckman and Vytlacil (2007b, Assumption B-2a, p. 5006): the indicator variable ${\rm 1\kern-3.99994pt1}(Z=t)$ can be used as the $Z^{[t]}$ in their assumption. Heckman and Vytlacil use their Assumption B-2a to identify the effect of the preferred treatment $t$ relative to the next-best treatment. Their complier group consists of those individuals who choose treatment $t$ under $Z=z$ and another treatment under $Z=z^{\prime}$. This can be a very heterogeneous group, as our examples will show. To paraphrase Heckman and Vytlacil (2007b, p. 5013): the mean effect of treatment $t$ versus the next best option is a weighted average over $t^{\prime}\in\mathcal{T}\setminus\{t\}$ of the effect of treatment $t$ versus treatment $t^{\prime}$, conditional on $t^{\prime}$ being the next best option, weighted by the probability that $t^{\prime}$ is the next best option. In contrast, we seek a complete characterization of all treatment effects that can be identified under this set of assumptions.

### v6

@@ 2 Targeting
It is important to note here that the values $U_{z}(t)$ and therefore the targeting maps $Z^{\ast}$ and $T^{\ast}$ are not observable; any assumption on targeting instruments and targeted treatments must be a priori and context-dependent. These prior assumptions have consequences that can be tested, however. As a simple example, $z=1$ targets $t=1$ in the $2\times 2$ model under monotonicity; this implies that $P(1|1)\geq P(1|0)$. We will show that targeting always yields linear inequalities on generalized propensity scores in the class of models defined by Assumptions 1 and 2.
2.1 Top targeted treatments and top alternative treatments
Definition 3 (Top targeted and top alternative treatments).
Take any observation $i$ in the population, and an instrument value $z\in\mathcal{Z}$.
  1. (i)
    If $z$ is a targeting instrument (that is, $z\in\mathcal{Z}^{\ast}$), let
    $V^{\ast}_{i}(z)=\max_{t\in\mathcal{T}^{\ast}(z)}(\bar{U}(t)+u_{it})$
    and $t^{\ast}_{i}(z)$ denote the maximizer, which is unique almost surely. We call $t^{\ast}_{i}(z)$ the top targeted treatment for observation $i$ under instrument value $z$.
    If $z\not\in\mathcal{Z}^{\ast}$, we define $V^{\ast}_{i}(z)=$-$\infty$.
  2. (ii)
    Also define
    $\underline{V}_{i}(z)=\max_{t\in\mathcal{T}\setminus\mathcal{T}^{\ast}(z)}\left(U_{z}(t)+u_{it}\right)$
    and let $\underline{t}_{i}(z)$ denote the almost surely unique maximizer. We call $\underline{t}_{i}(z)$ the top alternative treatment for observation $i$ under $z$.
By definition, the treatment $T_{i}(z)$ assigned to $i$ under $z$ can only be the top targeted treatment $t^{\ast}_{i}(z)$ (if $V^{\ast}_{i}(z)>\underline{V}_{i}(z)$) or the top alternative treatment $\underline{t}_{i}(z)$ (if $V^{\ast}_{i}(z)<\underline{V}_{i}(z)$). If $z$ is not a targeting instrument, then $T_{i}(z)$ must be $\underline{t}_{i}(z)$.
Definition 3 could be extended to broader settings than our ARUM model. On the other hand, the ARUM structure interacts with our definition of targeting to constrain treatment assignment in interesting ways. Suppose that an instrument value $z$ targets a treatment value $t$: $U_{z}(t)=\bar{U}(t)$. If some observation $i$ is assigned a treatment value $T_{i}(z)=t^{\prime}\neq t$, we must have
$\bar{U}(t)+u_{it}<U_{z}(t^{\prime})+u_{it^{\prime}}.$
Now suppose that an instrument value $z^{\prime}$ targets this $t^{\prime}$; then $U_{z}(t^{\prime})\leq U_{z^{\prime}}(t^{\prime})=\bar{U}(t^{\prime})$, so that
(2.1) $U_{z^{\prime}}(t)+u_{it}\leq\bar{U}(t)+u_{it}<U_{z}(t^{\prime})+u_{it^{\prime}}\leq U_{z^{\prime}}(t^{\prime})+u_{it^{\prime}}$
and $T_{i}(z^{\prime})\neq t$. If $t^{\prime}=0$, we can use Assumption 2 to get a stronger implication: since $U_{z}(0)=U_{z^{\prime\prime}}(0)=0$ for all $z^{\prime\prime}$, (2.1) holds for all $z^{\prime\prime}$ and $T_{i}(z^{\prime\prime})\neq t$. We summarize this in Proposition˜1.
Proposition 1 (Consequences of Targeting).
If $z$ targets $t$ and $T_{i}(z)=t^{\prime}\neq t$, then under Assumptions 1, and 2:
  1. (i)
    for all $z^{\prime}$ that target $t^{\prime}$, we have $T_{i}(z^{\prime})\neq t$;
  2. (ii)
    if $t^{\prime}=0$, then $T_{i}(z^{\prime\prime})\neq t$ for all $z^{\prime\prime}\in\mathcal{Z}$.
Proposition˜1 (i) implies that the events $T_{i}(z)=t^{\prime}$ and $T_{i}(z^{\prime})=t$ are mutually exclusive. Hence, the sum of their probabilities cannot exceed one:
(2.2) $P(t^{\prime}\mid z)+P(t\mid z^{\prime})\leq 1\quad\text{if }\;z\text{ targets }t\text{ and }z^{\prime}\text{ targets }t^{\prime}\neq t.$
Similarly, Proposition˜1 (ii) implies that the event $T_{i}(z)=0$ is disjoint from the event that $T_{i}(z^{\prime\prime})=t$ for any $z^{\prime\prime}\in\mathcal{Z}$. Since the probability of this latter event is at least $\max_{z^{\prime\prime}}P(t\mid z^{\prime\prime})$, we obtain the following implication:
(2.3) $P(0\mid z)+\max_{z^{\prime\prime}\in\mathcal{Z},z^{\prime\prime}\neq z}P(t\mid z^{\prime\prime})\leq 1\quad\text{if }\;z\text{ targets }t.$
2.2 One-to-one Targeting
@@
Under one-to-one targeting, we will often write $z=t$” if $z$ targets $t$; this is without loss of generality. Let us illustrate these varieties of targeting on Example 2.
@@
One-to-one targeting implies some useful restrictions on response-groups. Take an observation $i$ and $t\in\mathcal{T}^{\ast}$. Under one-to-one targeting, treatment $t$ is targeted only by $z=t$. As a result, by Proposition˜1, if $T_{i}(t)=t^{\prime}$ and $t^{\prime}\in\mathcal{T}^{\ast}\setminus\{t\}$ then $T_{i}(t^{\prime})\neq t$; and if $T_{i}(t)=0$ then $T_{i}(z)$ can never equal $t$. These restrictions are summarized in the following proposition.
@@
Under Assumptions 1, 2, and one-to-one targeting, take a targeted treatment $t\in\mathcal{T}^{\ast}$.
@@
    All response-groups $C_{R}$ with $R(t)=t^{\prime}\in\mathcal{T}^{\ast}\setminus\{t\}$ and $R(t^{\prime})=t$ are empty.
@@
    All response-groups $C_{R}$ with $R(t)=0$ and $R(z)=t$ for some $z\neq t$ are empty.
@@
This is easily rephrased in terms of the response-vectors of definition 1. First note that in the $3\times 3$ case, there are a priori $3^{3}=27$ response-vectors, $R=000$ to $R=222$, with corresponding response-groups $C_{000}$ to $C_{222}$. Groups $C_{ddd}$ are “always-takers” [^14] Observations in group $C_{000}$ are usually called the “never-takers”. We prefer not to break the symmetry in our notation. We hope this will not cause confusion. of treatment value $d$. All other groups are “compliers” of some kind, in that their treatment changes under some changes in the instrument. We will also pay special attention to some non-elemental groups. For instance, $C_{0\ast 2}$ will denote the group who is assigned treatment 0 under $z=0$ and treatment 2 under $z=2$, and any treatment under $z=1$. That is,
@@
One-to-one targeting implies the emptiness of four composite groups out of the 27 possible. For any treatment value $\tau$, Proposition 2(ii) rules out group $C_{10\tau}$ since this group has $R(1)=0$ and $R(0)=1$. It rules out $C_{\tau 01}$ as $R(1)=0$ and $R(2)=1$. This eliminates the composite groups $C_{10\ast}$ and $C_{\ast 01}$. The same argument applies to composite groups $C_{\ast 20}$ and $C_{2\ast 0}$, which have $R(2)=0$ and $R(1)=2$ or $R(2)=2$.
These four composite groups correspond to 10 elemental groups [^15] Specifically, they are: $C_{100},C_{101},C_{102},C_{001},C_{201},C_{020},C_{120},C_{220},C_{200}$, and $C_{210}$.. This still leaves us with 17 elemental groups, and potentially complex assignment patterns. Consider for instance Figure 2. It shows one possible configuration for the $3\times 3$ model; the positions for $P_{0},P_{1}$ and $P_{2}$ are consistent with one-to-one targeting.
$C_{222}$$C_{111}$$C_{110}$$C_{010}$$C_{000}$$C_{202}$$C_{002}$$C_{012}$$C_{112}$$C_{212}$$u_{i1}-u_{i0}$$u_{i2}-u_{i0}$$P_{0}$$P_{1}$$P_{2}$
@@
Heckman and Pinto (2018, pp. 16–20), Pinto (2022), and Kirkeboen, Leuven, and Mogstad (2016) also studied the $3\times 3$ model; they proposed sets of assumptions that identify some treatment effects. The example in Heckman and Pinto (2018, pp. 16–20) is rather specific. We show in Appendix˜F how to apply our framework to the Moving to Opportunity experiment studied in Pinto (2022). The setup in Kirkeboen, Leuven, and Mogstad (2016) is most similar to ours; we will return to the differences between our approach and theirs in Section 3.3.
@@ 2.3 Strict Targeting
Suppose for instance that the data comes from a randomized experiment, where the instrument value $z=t$ targets treatment $t$. If compliance is imperfect, an individual will trade off the benefits from switching to a treatment $t^{\prime}\neq t$ with the costs of the effort required. Strict targeting obtains when the cost of switching to $t^{\prime}$ do not depend on the value of $t$.
@@
2.3.1 Consequences of Strict Targeting
@@
Assumption 3 (Full strict targeting).
@@
Under full strict targeting, the values of $U_{z}(t)$ are given in Table 2. The functions $\underline{V}_{i}$ and $\underline{t}_{i}$ that define the top alternative treatment for a given observation are constant over $\mathcal{Z}\setminus\mathcal{Z}^{\ast}$. The assigned treatment $T_{i}(z)$ is $\underline{t}_{i}$ for all such values of $z$, as well as for $z\in\mathcal{Z}^{\ast}$ if
$V^{\ast}_{i}(z)=\max_{t\in\mathcal{T}^{\ast}(z)}(\bar{U}_{t}+u_{it})<\underline{V}_{i}=\max_{t\not\in\mathcal{T}^{\ast}(z)}(\underline{U}_{t}+u_{it});$
it is the maximizer $t^{\ast}_{i}(z)$ of $V^{\ast}_{i}(z)$ otherwise.
@@
$z\not\in Z^{\ast}$ - $\underline{U}(t)$
Table 2: Values of $U_{z}(t)$ under full strict targeting
Note that in a sense, all instrument values in $\mathcal{Z}\,\setminus\,\mathcal{Z}^{\ast}$ are equivalent under universal strict targeting. If $z$ and $z^{\prime}$ are two such values, then both functions $U_{z}$ and $U_{z^{\prime}}$ equal $\underline{U}$ on all of $\mathcal{T}$ and the counterfactual treatments $T_{i}(z)$ and $T_{i}(z^{\prime})$ must be $\underline{t}_{i}$ for any observation $i$.
2.4 Universal targeting
We now impose one-to-one targeting as well as strict targeting for every targeted treatment value:
Assumption 4 (Universal targeting).
If $t$ is in $\mathcal{T}^{\ast}$, then $t$ is strictly targeted, and it is uniquely targeted by $z=t$.
Note that Assumption˜4 implies Assumption˜3.
$z\in Z^{\ast}$ $t=z$ $t\neq z$
$\bar{U}(t)$ $\underline{U}(t)$
$z\not\in Z^{\ast}$ $\underline{U}(0)$, $\underline{U}(1)$, …, $\underline{U}(|\mathcal{T}|-1)$
Table 3: Values of $U_{z}(t)$ under universal targeting
Under Assumption˜4, Table˜2 becomes Table˜3 and we have:
Proposition 3 (Treatment assignment under universal targeting).
Take any observation $i$. Let $A_{i}$ be the (possibly empty) subset of $\mathcal{T}^{\ast}$ over which $T_{i}(t)=t$. Then under Assumptions 1 and 4,
@@ 2.4 Universal targeting
    if $z\in\mathcal{Z}\,\setminus\,A_{i}$ then $T_{i}(z)=\underline{t}_{i}$.
@@
    $\underline{t}_{i}$ belongs to the set $\bar{A}_{i}\equiv A_{i}\operatorname*{{\bigcup}}(\mathcal{T}\setminus\mathcal{T}^{\ast})$.
@@
    The pair $(A_{i},\underline{t}_{i})$ defines an elemental response group which we denote $C(A_{i},\underline{t}_{i})$. The family of sets $\{C(A,t)\;|\;A\subset\mathcal{T}^{\ast},t\in\bar{A}\}$ form a partition of the set of observations.
@@
If the set $A_{i}$ is non-empty and $\underline{t}_{i}\in A_{i}$, the observation $i$ is what one could call a strict $A_{i}$-complier: when the value of the instrument moves from $\mathcal{Z}\,\setminus\,A_{i}$ to a $t\in A_{i}$, observation $i$ switches from its top alternative treatment $\underline{t}_{i}$ to the top targeted treatment $t$. In the 3-by-3 model with $\mathcal{T}^{\ast}=\{1,2\}$, there are three groups of strict compliers: $C_{010}=C(\{1\},0)$, $C_{002}=C(\{2\},0)$, and $C_{012}=C(\{1,2\},0)$.
Universal targeting brings us very close to the main identifying assumption in Heckman and Vytlacil (2007b, Assumption B-2a, p. 5006): the indicator variable ${\rm 1\kern-3.99994pt1}(Z=t)$ can be used as the $Z^{[t]}$ in their assumption. Heckman and Vytlacil use their Assumption B-2a to identify the effect of the preferred treatment $t$ relative to the next-best treatment. Their complier group consists of those individuals who choose treatment $t$ under $Z=z$ and another treatment under $Z=z^{\prime}$. This can be a very heterogeneous group, as our examples will show. To paraphrase Heckman and Vytlacil (2007b, p. 5013): the mean effect of treatment $t$ versus the next best option is a weighted average over $t^{\prime}\in\mathcal{T}\setminus\{t\}$ of the effect of treatment $t$ versus treatment $t^{\prime}$, conditional on $t^{\prime}$ being the next best option, weighted by the probability that $t^{\prime}$ is the next best option. In contrast, we seek a complete characterization of all treatment effects that can be identified under this set of assumptions.
$t=0$ $t\in\mathcal{T}\setminus(\mathcal{T}^{\ast}\operatorname*{{\bigcup}}\{0\})$ $t\in\mathcal{T}^{\ast}$ $t^{\prime}=z^{\prime}$
$z\not\in\mathcal{Z}^{\ast}$ 0 $\underline{U}_{t}$ $\underline{U}_{t}$ $\underline{U}_{t^{\prime}}$
$z^{\prime}\in\mathcal{Z}^{\ast}\setminus\{t\}$ 0 $\underline{U}_{t}$ $\underline{U}_{t}$ $\bar{U}_{t^{\prime}}$
$z=t\in\mathcal{Z}^{\ast}$ 0 $\underline{U}_{t}$ $\bar{U}_{t}$ $\underline{U}_{t^{\prime}}$
Table 4: Universal targeting
It is easy to show that universal targeting yields two sets of testable implications. First, consider any treatment value $t\in\mathcal{T}$. It is clear from the first row of Table 4 that the mean utilities $U_{z}(t)$ are the same for all $z\in\mathcal{Z}\setminus\mathcal{Z}^{\ast}$. Since $0\not\in\mathcal{Z}^{\ast}$, it follows that
(2.4) $P(t|z)=P(t|0)\text{ for every }t\in\mathcal{T}\text{ and }z\in\mathcal{Z}\setminus\mathcal{Z}^{\ast}.$
Next, let $t\in\mathcal{T}^{\ast}$ be a targeted treatment value and take an instrument $z^{\prime}$ that targets a different treatment $t^{\prime}\neq t$. Comparing the rows of Table 4 reveals the ordering of utilities. When $z=t$, treatment $t$ is boosted; when $z=0$, it is at baseline; when $z=t^{\prime}$, the rival is boosted (drawing share away from $t$). This implies:
(2.5) $P(t|t)>P(t|0)>P(t|t^{\prime})\quad\text{for all }(t,t^{\prime})\in(\mathcal{T}^{\ast})^{2}\text{ with }t\neq t^{\prime}.$
Combining these inequalities implies that for all targeted instrument values $t\in\mathcal{T}^{\ast}$, the choice probability function $z\mapsto P(t\mid z)$ is strictly maximized at $z=t$.
Moreover, the “encouragement design” assumption of a recent paper by Bai, Huang, and Tabord-Meehan (2025) is exactly equivalent to universal targeting when (in their notation) $J_{0}>0$ and $z=0$ is what they call a “base state”. They show that encouragement design with a base state generates the following testable implications:
(2.6) $P(t|z)\leq P(t|0)\text{ if }z\text{ does not target }t.$
and that these implications are sharp: any set of propensity scores that satisfies these inequalities can be rationalized by an ARUM of encouragement design with a base state.
In our framework, combining (2.4) for $z\in\mathcal{Z}\setminus\mathcal{Z}^{\ast}$ and the rightmost inequality in (2.5) (strict inequality for rival-targeting $z$) yields (2.6). Conversely, the fact that choice probabilities must sum to one ensures that if $P(t\mid z)$ decreases for all non-target treatments, $P(t\mid t)$ must increase, satisfying the leftmost inequality of (2.5). Thus, under our assumptions, conditions (2.4) and (2.5) exhaust the testable implications of universal targeting.